Derive the expression of the trajectory when a charged particle of 

charge q is subjected to a magnetic field perpendicular to its plane. 

From here comment on the trajectory of the particle.


1
Expert's answer
2021-07-01T15:30:19-0400

Force that acts on the particle:


"F=qvB."

Since this force is perpendicular to the velocity vector, the particle experiences centripetal acceleration, which determines the net force:


"F=m\frac{v^2}{R}."

Equiate:

"qvB=m\frac{v^2}{R},\\\space\\
qB=\frac{mv}{R}.\\\space\\
R=\frac{mv}{qB}"

The trajectory is a circle with radius R.